40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6914     Entry time: Wed Jul 4 21:11:53 2012
Author: yuta 
Type: Update 
Category: Locking 
Subject: FPMI in vacuum is back 

I aligned FPMI and greens. There's no recognizable difference between before and after the vent.

What I did:
  1. Aligned Y arm to maximize Y green transmission.
  2. Used PZT1/2 to maximize TRY. But since PZT1 doesn't work so much, I had to align Y arm, too (mostly ETMY). This decreases green transmission, but I will leave it.
  3. Aligned BS and X arm to maximize TRX
  4. Fine tune them to minimize ASDC during FPMI lock, without decreasing TRX
  5. Aligned X end green to get TEM00 transmission.

Now, TRY and TRX are both  ~0.89.
Green transmission from Y and X arm are ~123 uW and ~275 uW respectively. Their max we got so far was ~200 uW and ~255 uW.
I still see clipped beam at AS, which I think is from the Faraday edge, as we found in elog #6897.
Below is the Sensoray capture of some ports, and MEDM screen shots to compare with before vent(see #6893).
There are two AS captures, one is without MI lock and the other is with MI lock. Note that PRM/SRM is misalined.

ALL_1025495266.pngMEDMscreenshotswithCOW_20120704.png


Next:
 - I will check ALS
 - I keep Y end green optics untouched since elog #6776, to use it as a reference. We need to realign them if tip-tilts are installed in vacuum, or PZTs are installed in both ends.

ELOG V3.1.3-